OpenStudy (anonymous):

Fe2O3 + 3CO→ 2Fe + 3CO2 A. How many moles of CO2 result when 50.0 moles of Fe2O3 react? B. How many grams of Fe result when 125grams of Fe2O3 react ?

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) N of Fe2O3=50/159.69=0.3131 moles of Fe2O3 look for the Mole ratie 1:3:2:2:3 Now compare the mole ratio between Fe2O3 and CO2 1:3 0.3131:x 0.3131*3/1=x mole x=0.939 moles Now if we used 125 grams of Fe2O3 125/159.69=0.7827 moles of Fe2O3 look for the mole ratio of iron oxide to iron 1:2 so the moles of iron is 1.5655 moles of iron Number of Moles=mass/molar mass 1.5655*55.8=87.3 grams of iron